What is the standard form of an equation that passes through the points,
(14, 8) and (3, 11) ?

Answer by ilana(307)

The standard form of a line is written in the form Ax+By+C=0.
With two points we can find their slope, then write the line in point-slope form, and then algebraically change the form.
The slope is (11-8)/(3-14)=-3/11.
So the line in point-slope form is y-11=-(3/11)(x-3).
Multiplying the equation by 11, 11y-121=-3x+9.
So this line is 3x+11y-112=0.